# On the Computation of BD-Rate over a Set of Videos for Fair Assessment of Performance of Learned Video Codecs

### v1

@@ Abstract
The Bjøntegaard Delta (BD) measure is widely employed to evaluate and quantify the variations in the rate-distortion (RD) performance across different codecs. Many researchers report the average BD value over multiple videos within a dataset for different codecs. We claim that the current practice in the learned video compression community of computing the average BD value over a dataset based on the average RD curve of multiple videos can lead to misleading conclusions. We show both by analysis of a simplistic case of linear RD curves and experimental results with two recent learned video codecs that averaging RD curves can lead to a single video to disproportionately influence the average BD value especially when the operating bitrate range of different codecs do not exactly match. Instead, we advocate for calculating the BD measure per-video basis, as commonly done by the traditional video compression community, followed by averaging the individual BD values over videos, to provide a fair comparison of learned video codecs. Our experimental results demonstrate that the comparison of two recent learned video codecs is affected by how we evaluate the average BD measure.
Index Terms—  rate-distortion performance over a dataset, average bjøntegaard delta metric, learned video compression.
@@ 1 Introduction
Comparing performances of different video codecs is a critical yet challenging task. A widely used measure for this purpose is the Bjøntegaard Delta (BD)[1] measure, which calculates the average difference between the rate-distortion (RD) curves of codecs, where the distortion is typically quantified by the peak signal-to noise ratio (PSNR) [2]).
In the classic video compression community, the RD curve and BD metric is typically computed per video. When we compare different codecs over a dataset, we compute the average of the BD rates over individual videos in the dataset. However, in the learned video compression community, a common practice among researchers is to average the RD curves over all videos in a dataset before calculating the BD metric. This approach, while seemingly straightforward, can lead to significant inaccuracies. When RD curves are averaged across videos, even a single video can skew the average RD curve, leading to misleading average BD values. Consequently, a model that performs well on most videos might appear less effective when evaluated on the average RD curve. This issue becomes more prominent when the operating bitrate range of different codecs do not exactly match as we show by the analysis of a simple case in Section 3.
A more accurate and fair assessment can be achieved by calculating the RD curve and the BD metric for each video individually and then averaging BD values, which is a common practice in the classic video compression community. This method ensures that the performance differences are measured consistently across all videos, providing a more reliable comparison of codecs. Furthermore, reporting individual video results avoids the pitfalls associated with averaging RD curves, offering a clearer insight into codec performance.
In the following, after a brief overview of related work in Section 2, we analyze the impact of averaging RD curves on the BD metric in a simplistic case in Section 3. We then provide experimental results to compare two recent learned video codecs in Section 4 showing how averaging RD curves can mislead performance comparisons. We hope that this paper helps improving evaluation practices in the learned video compression community, promoting more accurate and fair comparison of different video compression models by averaging per-video BD values instead of averaging RD curves.
@@ 2 Related work and Contributions
Traditional video codecs, such as those described by the High Efficiency Video Coding (HEVC) [3] and the Versatile Video Coding (VVC) [4] standards, have long been the cornerstone of video compression technology. These conventional codecs employ human-engineered algorithms and heuristics to achieve efficient compression by leveraging sophisticated temporal and spatial prediction methods.
In contrast, recent advancements in deep learning lead to learned video compression, which learn the best transform and statistics of the latent variables from large video collections by end-to-end optimization of neural networks. These models have successfully integrated motion compensation and temporal context exploitation, significantly enhancing compression performance[5, 6, 7, 8]. Researchers have introduced innovative concepts, such as conditional coding[9] and advanced entropy modeling[10, 11, 12, 13] to further optimize compression efficiency, while others have explored bi-directional coding strategies that leverage both forward and backward motion compensation[14, 15, 16, 17, 18, 19]. The continuous evolution of these methods reflects a growing interest in achieving greater adaptability and performance, ultimately aiming to surpass the limitations of traditional video coding standards.
Both communities compare performances of different codecs using BD values, but they employ different practices in the computation of BD values over a dataset. Notably, the traditional video coding community compute RD curves and BD values per video and then average the BD values, while the learned video compression community often compute the average of RD curves over the dataset, and then compute a BD value based on the average RD curve.
The main contribution of this paper is to show that computing the average of RD curves is not a reliable method to compare video codecs, and the learned video compression community should also compute RD curves and BD values per video and then average BD values for a fair comparison of learned video codecs. More specifically,
  • We provide an analysis of a simplistic scenario of two codecs with linear RD curves in Section 3 to show that averaging RD curves yield inconsistent results when the bitrate ranges of two codecs do not exactly align.
  • We provide experimental results in Section 4 to show that computing RD curves and BD values per video and then averaging BD values results in more reliable and fair comparison of two recent learned video codecs.
3 Analysis for a Simple Scenario
@@
Hence, the slope is $m=(\Delta P_{1}+\Delta P_{2})/(\Delta B_{1}+\Delta B_{2})$.
@@
Again, the slope is $m=(\Delta P_{1}+\Delta P_{2})/(\Delta B_{1}+\Delta B_{2})$.
In order to obtain the same BD-rate performance for both codecs, the average RD curves of two codecs must overlap. Therefore, the equation of the lines should be the same. So, if we substitute the first points of the RD curves into the equations for each codec, we have
$\dfrac{P_{1}^{1}+P_{2}^{1}}{2}$ $=\dfrac{\Delta P_{1}+\Delta P_{2}}{\Delta B_{1}+\Delta B_{2}}\cdot\dfrac{R_{1}^{1}+R_{2}^{1}}{2}+b_{1}$
which implies
$b_{1}=\dfrac{P_{1}^{1}+P_{2}^{1}}{2}-\dfrac{\Delta P_{1}+\Delta P_{2}}{\Delta B_{1}+\Delta B_{2}}\cdot\dfrac{R_{1}^{1}+R_{2}^{1}}{2},$ (1)
and
$\dfrac{P_{1}^{1}+P_{2}^{1}+\Delta P_{2}}{2}$ $=\dfrac{\Delta P_{1}+\Delta P_{2}}{\Delta B_{1}+\Delta B_{2}}\cdot\dfrac{R_{1}^{1}+R_{2}^{1}+\Delta B_{2}}{2}+b_{2}$
which implies
$b_{2}=\dfrac{P_{1}^{1}+P_{2}^{1}+\Delta P_{2}}{2}-\dfrac{\Delta P_{1}+\Delta P_{2}}{\Delta B_{1}+\Delta B_{2}}\cdot\dfrac{R_{1}^{1}+R_{2}^{1}+\Delta B_{2}}{2}.$ (2)
If we equate $b_{1}$ and $b_{2}$, we have
$\Delta P_{2}\cdot\Delta B_{1}=\Delta P_{1}\cdot\Delta B_{2}.$ (3)
Even in this simple case, if a certain condition as in Eq. (3) is not satisfied, then the average of BD-rate for individual videos is not equal to BD-rate calculated from the average RD curves.
@@
Fig. 2: Rate-Distortion curves for the seven (7) videos in the UVG[20] dataset. The last graph is the average RD curve. Observe that ReadySetGo is the only video where the average RD curve for DCVC-DC[13] is on top of that of Yılmaz [19].
Table 1: BD-BR values for Yılmaz [19] in reference to DCVC-DC[13]. Observe that computing average of per-video BD-BR values vs. BD-BR from the average RD curve lead to conflicting comparative performances.
@@
4 Experiments: Comparison of
Two Learned Video Codecs
@@ 4.1 Experimental Setup
We evaluate computing the BD rate over a dataset by averaging per-video BD rates vs. BD rate from the average RD curve to compare two recent learned video compression models: DCVC-DC[13], which performs sequential coding and learned bidirectional coding[19]. We have chosen these two models because, as shown by our results, one model [19] clearly outperforms the other by averaging per-video BD rates, while the average RD curve suggests otherwise.
Test Dataset We benchmarked the models using the UVG dataset[20], which contains seven videos with different characteristics at 1920 $\times$ 1080 resolution and is commonly used for benchmarking video codecs.
@@
Fig. 3: Average RD curve for the UVG dataset if ReadySetGo is excluded. The average RD curve for DCVC-DC[13] is on top of that of [19], although we removed the only video where the curve for DCVC-DC[13] is on top of that of [19].
@@
Fig. 4: Average RD curve for the UVG dataset if Beauty is excluded. Excluding Beauty, where Yilmaz_2024 is superior, puts the RD curve for Yilmaz_2024 over that of DCVC-DC [13], which is inconsistent with other average RD curves.
4.2 Results
Fig. 2 presents RD curves plotted for four different RD points across two different models, where x and y axes represent bits-per-pixel (bpp) and the PSNR, respectively. The first seven figures display per-video results, while the final figure shows the averaged plot over the UVG dataset, using the four-point averages, as commonly presented in the literature. In these results, the %BD-BR in the headings indicates the percentage rate gain of the Yilmaz_ICIP2024[19] model compared to the Li_CVPR2023[13] model. It is important to note that each video exhibits different bitrate (bpp) and PSNR ranges. For instance, the Beauty sequence, which contains complex and challenging motion patterns, tends to saturate quickly in quality, with limited PSNR improvements at high bitrates. In fact, the maximum PSNR value for the Beauty sequence can be lower than the minimum PSNR value when observed on the average RD curve, despite the sequence having much higher bitrates. This Beauty sequence effectively pulls down the average RD curve. On the other hand, sequences with less motion, such as Bosphorus and Honeybee, are able to achieve high PSNR values even at lower bitrates, and thus have a positive, upward-shifting effect on the average RD curve.
We observe in Fig.2 that Yilmaz_ICIP2024 outperforms Li_CVPR2023 for all sequences except ReadySetGo. When looking at the average Bjøntegaard Delta bitrate (BD-BR) gain across all videos, Yilmaz_ICIP2024 model shows a superior performance of -7.03%. However, if we calculate the BD metric based on the average bpp and PSNR values without accounting for the per-video differences, we may mistakenly conclude that the Yilmaz_ICIP2024 model is 3.56% inferior, as observed in the average RD curve plot. The reason for Yilmaz_ICIP2024 appearing inferior at high bitrates on the average RD curve is twofold: the downward-pull effect of the Beauty sequence, and the fact that the Li_CVPR2023 method has a higher endpoint at 0.1 bpp, 34.75 dB, compared to the Yilmaz_ICIP2024 method endpoint at 0.2 bpp, 35.50 dB.
In order to show that computing the average of RD curves leads to inconsistent results, we present Figs. 3 and 4, where the average RD curve is calculated after excluding one sequence from the video set. In Fig. 3, we exclude ReadySetGo sequence, which is the only sequence where Yilmaz_ICIP2024 model is inferior to Li_CVPR2023 model. Despite Yilmaz_ICIP2024 model demonstrates superior performance over Li_CVPR2023 model for each individual sequence used to compute the average curve, the average RD curve for Li_CVPR2023 appears on top at high bitrates, which is clearly an inconsistent result. In Fig. 4, we exclude Beauty sequence. Despite the fact that Yilmaz_2024 model is superior in that sequence, exclusion of the Beauty sequence, which had a downward pull on the average curve, reveals that the Yilmaz_2024 model, which initially appeared inferior in terms of BD-BR (3.56%) when all sequences were included, actually demonstrates superior BD-BR (-1.74%) after exclusion. This result also demonstrates potential misleading nature of averaging RD curves, as it highlights how certain sequences can skew overall performance assessment.
@@ 5 Conclusion
This paper shows how using average RD curves can be misleading for the computation of BD-rate over a test set when comparing learned video compression models. To illustrate this more concretely, we analyzed a hypothetical case for two codecs with linear RD curves as well as conducted experiments to compare averaging per-video BD values vs. computing a BD value from the average RD curve for two recent learned codec models on the UVG dataset. Our results revealed that a model that appears inferior when looking at the average RD curve can actually be superior when the per-video results are computed and then averaged. We showed both in the analysis of a simplistic setting and in the experiments with real codecs that averaging RD curves can yield inconsistent and unreliable results. Therefore, we conclude that averaging per-video BD values provides a more fair comparison of learned video codecs, while the performance derived from average RD curves may be misleading. We argue that researchers should carefully consider the findings in this paper as the current practice in the learned video compression community to compute BD rate over a set of videos can lead to misleading conclusions about the relative strengths of different learned video compression models.

### v2

# The Practice of Averaging Rate-Distortion Curves over Testsets to Compare Learned Video Codecs Can Cause Misleading Conclusions

@@ Abstract
This paper aims to demonstrate how the prevalent practice in the learned video compression community of averaging rate-distortion (RD) curves across a test video set can lead to misleading conclusions in evaluating codec performance. Through analytical analysis of a simple case and experimental results with two recent learned video codecs, we show how averaged RD curves can mislead comparative evaluation of different codecs, particularly when videos in a dataset have varying characteristics and operating ranges. We illustrate how a single video with distinct RD characteristics from the rest of the test set can disproportionately influence the average RD curve, potentially overshadowing a codec’s superior performance across most individual sequences. Using two recent learned video codecs on the UVG dataset as a case study, we demonstrate computing performance metrics, such as the BD rate, from the average RD curve suggests conclusions that contradict those reached from calculating the average of per-sequence metrics. Hence, we argue that the learned video compression community should also report per-sequence RD curves and performance metrics for a test set should be computed from the average of per-sequence metrics, similar to the established practice in traditional video coding, to ensure fair and accurate codec comparisons.
Index Terms—  learned video compression, performance evaluation over test set, rate-distortion curve, average BD-rate
@@ 1 Introduction
Comparing performances of different video codecs is a critical yet challenging task in video compression research. While measuring performance through rate-distortion (RD) curves provides insights into codec behavior, the methodology of presenting and analyzing these curves can significantly impact conclusions. The Bjøntegaard Delta (BD) [1] measure, which calculates the average difference between RD curves, where distortion is typically quantified by peak signal-to-noise ratio (PSNR) [2], is commonly used to quantify these comparisons.
The classic video compression community evaluate codecs by computing RD curves and performance metrics, such as the BD rate, per-video, then averaging these metrics across a dataset. In contrast, the learned video compression community commonly present an RD curve averaged across all videos in a dataset, and then compute a single BD rate from this average RD curve. This latter practice, while appearing to be a convenient approach, can lead to misleading conclusions about codec performance. When RD curves are averaged across videos with varying characteristics, even a single video can dramatically skew the average curve. For instance, a video with distinct RD characteristics can disproportionately influence the average, potentially masking a codec’s superior performance across most sequences. This problem becomes particularly severe when different codecs operate at varying bitrate ranges - a common scenario in learned compression where models may adapt differently to content characteristics. To ensure fair and accurate codec assessment, we advocate for following the traditional video coding community’s approach: computing RD curves and performance metrics for each video individually before averaging.
In the following, Section 2 reviews related work and discusses contributions of this paper, Section 3 presents an analytical analysis of a simple case demonstrating how averaged RD curves can lead to misleading conclusions, and Section 4 offers experimental evidence comparing two recent learned video codecs, revealing how computing BD rate from averaged RD curves leads to a different conclusion compared to averaging per-sequence BD-rates. By highlighting the pitfalls of averaging RD curves and advocating for evaluation of average BD-rates as average of BD-rates calculated from per-sequence RD curves, we hope to influence the learned video compression community to adopt the more accurate and reliable assessment methodology that has been a common practice in the classic video coding community.
@@ 2 Related work and Contributions
Traditional video codecs, such as those described by the High Efficiency Video Coding (HEVC) [3] and the Versatile Video Coding (VVC) [4] standards, employ rigorous evaluation methodologies where codec performance is assessed on a per-sequence basis. These conventional codecs are compared using consistent metrics computed individually for each sequence, ensuring fair assessment across different content types. In contrast, the learned video compression community has widely adopted a different evaluation methodology. Despite significant advances in compression techniques through deep learning approaches [5, 6, 7, 8, 9, 10, 11, 12, 13], the evaluation practices have diverged from traditional practices.
Examination of recent literature reveals the prevalence of averaging RD curves across entire datasets before performance analysis. For instance, Chen and Peng [14] explicitly stated ”The average bpp and PSNR-RGB are computed across the entire dataset, treating it as a single input sequence.” This practice is evident in numerous influential works, including scale-space flow methods [5], hierarchical quality enhancement approaches [15], and neural representation-based compression [16, 17]. The issue extends beyond individual papers to affect the entire field. Recent works in learned bi-directional coding [18, 15, 19, 20, 21] predominantly rely on averaged RD curves for performance assessment. This practice has become so common that even some benchmark platforms [22] primarily present averaged curves rather than per-sequence analyses.
The main contribution of this paper is to demonstrate why the practice of averaging RD curves can lead to misleading conclusions about codec performance. Specifically:
  • We provide an analytical analysis in Section 3 showing how averaged RD curves can produce inconsistent results, particularly when codecs operate in different bitrate ranges - a common case in learned compression.
  • Through experiments in Section 4, we demonstrate how the practice of averaging RD curves can overshadow performance differences between codecs, using two recent learned compression methods as a case study.
We would like to emphasize that while the per-sequence computation of RD curves and BD rates is an established approach in the traditional video coding community, there is no published literature that shows the pitfalls of computing a single BD rate from the average RD curve, which is the de facto approach in the learned video compression community.
3 Why Averaging RD Curves Misleads Codec Assessment:
Analysis for a Simple Case
@@ 3 Why Averaging RD Curves Misleads Codec Assessment:
Both curves have slope $m=(\Delta P_{1}+\Delta P_{2})/(\Delta B_{1}+\Delta B_{2})$. For these curves to overlap and show equivalent performance, their y-intercepts must be equal. After substituting these points into $P=mR+b$ and solving for the intercepts, we find that they are equal only when:
$\Delta P_{2}\cdot\Delta B_{1}=\Delta P_{1}\cdot\Delta B_{2}$ (1)
Even in this simple case, if this condition is not satisfied, then the average RD curves will indicate different performance levels, contradicting the per-sequence evaluation results where the codecs perform identically.
@@
Fig. 2: Rate-Distortion curves for the seven (7) videos in the UVG[23] dataset. The last graph is the average RD curve. Observe that ReadySetGo is the only video where the average RD curve for Li_CVPR2023 is on top of that of Yilmaz_ICIP2024.
Table 1: BD-BR values for Yılmaz_ICIP2024 in reference to Li_CVPR2023. Observe that computing average of per-video BD-BR values vs. BD-BR from the average RD curve lead to conflicting comparative performances.
@@
4 Experimental Evidence
@@ 4.1 Experimental Setup
To demonstrate how averaging RD curves can lead to misleading conclusions about codec performance, we analyze two recent learned video compression models: Li_CVPR2023 (DCVC-DC)[13], which performs sequential coding, and learned bidirectional coding Yilmaz_ICIP2024[24]. These models provide an excellent case study, as one model shows superior performance when evaluated properly on a per-sequence basis, yet appears inferior when assessed using averaged RD curves. We conducted experiments using the UVG dataset[23], which contains seven videos with diverse characteristics at 1920 $\times$ 1080 resolution. This dataset is particularly suitable for our analysis as it includes sequences with varying complexity, motion patterns, and operating ranges - factors that can significantly impact how averaging affects performance assessment.
@@
Fig. 3: Average RD curve for the UVG dataset if ReadySetGo is excluded. The average RD curve for Li_CVPR2023 is on top of that of Yilmaz_ICIP2024, although we removed the only video where the curve for Li_CVPR2023 is on top of that of Yilmaz_ICIP2024.
@@
Fig. 4: Average RD curve for the UVG dataset if Beauty is excluded. Excluding Beauty, where Yilmaz_ICIP2024 is superior, puts the RD curve for Yilmaz_ICIP2024 over that of Li_CVPR2023, which is inconsistent with other average RD curves.
4.2 Analysis of RD Curve Averaging Effects
Fig. 2 presents per-sequence and averaged RD curves (PSNR vs. bits-per-pixel (bpp)) for both models. The per-sequence plots reveal that Yilmaz_ICIP2024 outperforms Li_CVPR2023 on all sequences except ReadySetGo. The performance differences vary significantly across sequences, with substantial gains on sequences like Bosphorus (-25.16% BD-BR) and Honeybee (-23.19% BD-BR), as shown in Table 1.
However, examining the averaged RD curve (last plot in Fig. 2) leads to a contradictory conclusion. Despite superior performance on most individual sequences, the averaged curve suggests Yilmaz_ICIP2024 is 3.56% inferior. This misleading result stems from two key factors:
1. Operating Range Differences: The Beauty sequence operates in a distinctly different quality range, with its maximum PSNR being lower than other sequences’ minimum values. When averaged, this sequence disproportionately pulls the overall curve down.
2. Sequence Characteristics: Videos like Beauty, with complex motion patterns, saturate quickly in quality with limited PSNR improvements at high bitrates. In contrast, simpler sequences like Bosphorus and Honeybee achieve high PSNR values even at lower bitrates, creating an upward-shifting effect on the average curve.
4.3 Impact of Individual Sequences
To further illustrate how averaging RD curves can mask true performance differences, we conducted two experiments:
First, we excluded ReadySetGo - the only sequence where Yilmaz_ICIP2024 shows inferior performance. As shown in Fig. 3, despite Yilmaz_ICIP2024 being superior on all remaining sequences, the averaged RD curve still shows Li_CVPR2023 outperforming at high bitrates - a clearly inconsistent result.
Second, we excluded the Beauty sequence (Fig. 4). Despite Yilmaz_ICIP2024 showing superior performance on this sequence, its exclusion dramatically changes the average RD curve comparison. The model that appeared inferior when all sequences were included (3.56% worse) now demonstrates superior performance (-1.74% BD-BR).
These results highlight how averaging RD curves can lead to conclusions that directly contradict per-sequence performance metrics. The practice of averaging RD curves, while common in the learned video compression community, fails to capture the true relative strengths of different codecs and can lead to misleading assessments of codec performance.
@@ 5 Conclusion
This paper exposes a methodological issue in the evaluation of learned video compression models: the widespread adopted practice of averaging RD curves across video sequences leads to potentially misleading performance assessments. Through both analytical analysis and real-world experiments, we demonstrate how this common practice can affect codec performance evaluation and even reverse comparative outcomes. Our results reveal that a codec showing superior performance on most individual sequences can paradoxically appear inferior when evaluated using average RD curves.
The implications of our findings extend beyond just BD-rate computation. Our experimental results demonstrate how sequence-specific characteristics and varying operating ranges can disproportionately influence averaged RD curves. This influence becomes particularly problematic when different codecs operate in varying bitrate ranges - a common scenario in learned compression where models may adapt differently to content characteristics.
Most importantly, we demonstrate that averaging per-video performance metrics, rather than averaging RD curves, is a more reliable and fair assessment methodology. We emphasize that while this is already an established approach in the traditional video coding community, there is no published literature that shows the pitfalls of computing a single BD rate from the average RD curve, which is the de facto approach in the learned video compression community.
As the field of learned video compression continues to advance rapidly, ensuring fair and accurate performance assessment becomes increasingly crucial. We encourage researchers to carefully consider these findings and adopt per-sequence evaluation practices, as misleading performance metrics could significantly impact research directions and algorithmic choices in this rapidly evolving field.